One number is 4 less than another number. The sum of the two numbers is four times their difference. What are the two numbers?
Answer by RAY100(1637)

Let n= smaller #,,,Let m=larger #
.
1) m-4 =n
.
2) n+m = 4(m-n)
.
from 2) n+m = 4m-4n
.
5n=3m
.
5/3 (n) = m
.
subst into (1)
.
{5/3(n)] -4 = n
.
2/3 n =4
.
n=6,,,,,,,,,from(1) m-4 = 6,,,,m=10
.
check,,,(1) 10-4=6,,,,,ok
,
(2) 6+10 =16 = 4(10-4) = 4*4,,,,,ok